OpenStudy (anonymous):

The volume of a hexagonal prism is 10 in3. What is the height of the prism if the base area is 5 in3? 50 in 2 in 50 in 2 in

OpenStudy (anonymous):

h prisms = V/S V-volume S -area h - height h = 10/5 = 2

OpenStudy (anonymous):

:) tnx a lot.

OpenStudy (anonymous):

:)